OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 3) and is perpendicular to the line y = –x – 3?

OpenStudy (anonymous):

y-3=-(x-2)

OpenStudy (anonymous):

these are my options: y = –x – 1 y = –x + 5 y = x + 1 y = x + 5

OpenStudy (anonymous):

which one would you choose?

OpenStudy (anonymous):

which one corresponds to my answer?

OpenStudy (anonymous):

the first one?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i really don't know